OpenStudy (anonymous):

40 foot ladder is leaning against a building so that the distance from the bottom of the ladder to the building is 8-ft more than the distance the ladder reaches up the building. Find how far up the building the ladder reaches.

OpenStudy (anonymous):



OpenStudy (anonymous):

Pythagoras

OpenStudy (anonymous):

x^2 + (x + 8)^2 = 40^2 x^2 + 8^2 + 16x + x^2 = 40^2

OpenStudy (anonymous):